Suppose a single person earns $50,700 per year and also has other monthly installment obligations of $550. The purchase of a home requires a 5% down payment and an escrow of 3.2% per year to cover property taxes and hazard insurance. Private mortgage insurance (PMI) of .78% of the loan value per year is also required. The loan underwriter will apply a DTI (debt-to-income) or “back ratio” of 43%. The 30-year monthly loan payment factor is .004774 for a loan at 4% interest per year. This person can afford a home priced at:

$521,669

$260,834

$169,500

$162,000
Answer #1
A DTI(Debt to Income Ratio) 43%
B Annual Income $50,700
C=A*B Total Monthly Debt Expense $21,801
Assume , Home Priced at =X
Down Payment =5%
Loan amount =0.95X
D Annual Loan Repayment =12*0.004774*0.95X 0.0544236 *X
E Escrow Amount (3.2%) 0.032 *X
F Private Mortgage Insurance =0.0078*0.95X 0.00741 *X
G=D+E+F Total Payment for the property 0.0938336 *X
Other Debt Obligation=550*12 $6,600
Total Monthly Debt Expense=0.093834X+6600= $21,801
0.093834X=21801-6600= $15,201
X=15201/0.093834= $162,000
ANSWER:
$162,000